If a man runs at a speed of 13kph, what would be his folded distance in 7 minutes?
1
Expert's answer
2020-08-19T12:56:43-0400

The distance traveled by man

"d=vt=\left(13\times \frac{1000\:\rm m}{3600\:\rm s}\right)\times (7\times 60\:\rm s)=1517\: m"